The enthalpy change For the reaction of 50 ml if ethylene with 50 ml of H at 1.5 atm pressure. /\ H= -0.31 kj. What is the /\ U??

Grade:11

3 Answers

Arun
6 years ago
Dear Mohit
 

the reaction can be written as follows:

C2H4 + H2 → C2H6

the ratio of reactant is 1:1 and product is also forming in ratio of 1,∴ 50 ml of ethylene on reaction with 50 ml of H2 will produce 50 ml. of ethane. change in volume is 50ml (ethane)-100ml (C2H4 + H2) = 50ml= 0.05L

P ΔV=1.5*0.05= 0.075Latm = 7.5joules

ΔU = ΔH + P ΔV = -0.31KJ + 7.5J = -.3025KJ

 

Regards

Arun (askIITians forum expert)

 

KAVITA YADAV
4 years ago
Sjshshjhjjhjdjhhjsjdvdhdhdbdjdbdjdbdidbksjshdbdkshsbdjfhfbdjdhdjdjdidjdjdnskshdhidehakpsnswlebdjdognjcdsheinxkhfkrnzlhidnldrkbskwjpdjwjdndkdbdjkdnddfjbfjchdodndnzxjdjebdkrhdjdueiejehdnkfdojensnfjridhfbkfkdnddkzks
/\U=0.3024kJ
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your

the reaction can be written as follows:
C2H4 + H2 → C2H6
the ratio of reactant is 1:1 and product is also forming in ratio of 1,∴ 50 ml of ethylene on reaction with 50 ml of H2 will produce 50 ml. of ethane. change in volume is 50ml (ethane)-100ml (C2H4 + H2) = 50ml= 0.05L
PΔV=1.5*0.05= 0.075Latm = 7.5joules
ΔU = ΔH + P ΔV = -0.31KJ + 7.5J = -.3025KJ

Thanks and Regards